OpenStudy (anonymous):

the sum of two consecutive intergers is 83 What are the Two numbers?

OpenStudy (anonymous):

x=#1 and x+1=#2

OpenStudy (anonymous):

2x=82

OpenStudy (anonymous):

(x) +(x+1) = 83 2x + 1 = 83 2x = 83-1 x = 82/2 x = 41 so the two numbers r 41 and 42

OpenStudy (anonymous):

tks guys !!!